A footwear factory sold 600 pairs of slippers. the number sold at market stalls is twice the number sold at department stores and the number sold by foot salesmen is one-sixth of the number sold at market stalls. how many pairs of slippers were sold at market stalls?
Answer by lwsshak3(11628)

A footwear factory sold 600 pairs of slippers. the number sold at market stalls is twice the number sold at department stores and the number sold by foot salesmen is one-sixth of the number sold at market stalls. how many pairs of slippers were sold at market stalls
***
let x=number of pairs of slippers sold at dept stores
2x=number of pairs of slippers sold at market stalls
2x/6=number of pairs of slippers sold by foot salesmen
x+2x+2x/6=600
x+2x+x/3=600
lcd:3
3x+6x+x=1800
9x=1800
x=200
2x=400
how many pairs of slippers were sold at market stalls? 400